OpenStudy (anonymous):

$1500 is invested for 6 months at an annual simple interest rate of 14%. How much interest will be earned?

OpenStudy (dumbcow):

Interest = P*i*t = 1500*0.14*(1/2) because 6 months is 1/2 a year =105

OpenStudy (anonymous):

so if you wanted to find the percent you would i*r*t?

OpenStudy (dumbcow):

whats r? if you want to find the interest rate as percent i% = (Interest/P*t) *100

OpenStudy (anonymous):

rate

OpenStudy (dumbcow):

ahh ok im using i for the rate,

OpenStudy (anonymous):

ohh so interest & rate are the same?

OpenStudy (dumbcow):

kinda i = interest rate when i typed out Interest i meant the dollar amount of interest earned so interest rate and rate are the same thing but Interest is different

OpenStudy (dumbcow):

sorry for confusing you

OpenStudy (anonymous):

a little lol

OpenStudy (anonymous):

but thanks for the mini lesson

OpenStudy (dumbcow):

:)